OpenStudy (anonymous):

Juan always saves the same amount from his weekly allowance. The table shows how much he has saved at different times. How much will Juan have saved in 14 wk?

OpenStudy (anonymous):

PLEASE HELP

Mehek (mehek14):

table?

OpenStudy (anonymous):

WILL GIVE MEDAL

OpenStudy (anonymous):

Juan always saves the same amount from his weekly allowance. The table shows how much he has saved at different times. How much will Juan have saved in 14 wk? $ Juan’s Savings Weeks 3 6 9 12 Amount saved 19 28 37 46

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

@MAli13chineta @Mchilds15

Mehek (mehek14):

ok so find the rate of change \(\dfrac{28 - 19}{6-3} = 3\)

OpenStudy (anonymous):

ok

Mehek (mehek14):

now find the slope 19 = 3*3 + b 10 = b so your equations is y = 3x + 10 sub 14 for x y = 3*14 + 10 y = 42 + 10 y = 52

OpenStudy (anonymous):

so answer is 52?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

yes

Mehek (mehek14):

yea